What is the intensity of sound waves produced by a piano at a distance of 1.5 m when the power output of the piano is 3.0 W? 


1
Expert's answer
2021-10-12T16:23:07-0400

The intensity of sound waves is subjected to the inverse square law, it is the sound power divided by area:


"I=\frac P{4\pi r^2}=0.106\text{ W/m}^2."